# Pull-Based Query Scheduling for Goal-Oriented Semantic Communication

### v1

@@ I Introduction
When analyzing long-term effectiveness, the axioms of classical expected utility theory (EUT), where the weight of an occurrence is determined by its probability, could be challenged in the context of risk-aware semantic decision-making. Decisions involving risky prospects, perceptual context evaluation, and semantic utility valuation reveal several effects the utility theory fails to capture. One prominent effect is certainty, which suggests that merely probable outcomes are undervalued compared to certain ones. For example, in human-centric applications, there is often a preference to guarantee achieving a significant portion of a goal rather than risk achieving the entire goal with less certainty. Decision-making involving human agents and human-in-the-loop control of critical applications tends to be more conservative, prioritizing safety and performance guarantees over risk-taking, even when the latter has the potential to maximize the utility function. Cumulative prospect theory (CPT) was developed as an alternative framework to EUT, addressing subjectivity and behavioral effects in decision-making [10, 11, 12]. This theory introduces a more sophisticated approach by linking utilities to gains and losses via a value function and transforming the probability distribution of outcomes into decision weights through a weighting function. This nuanced model better captures the complexities of decision-making under uncertainty and reflects how individuals semantically evaluate and prioritize outcomes in real-world scenarios.
@@
In this paper, we investigate a pull-based end-to-end status update system in which multiple sensing agents monitor a source and provide updates to a group of actuation agents, each taking actions and working toward achieving heterogeneous goals at the endpoint. The source comprises multiple attributes, and updates about these attributes are routed through a hub that dynamically determines which specific attribute to query at any given time. To assess the effectiveness of updates, we utilize the grade of effectiveness concept introduced in [13], incorporating two key semantic metrics: freshness and usefulness. Freshness quantifies how quickly an update becomes outdated after successfully reaching its endpoint. Usefulness, on the other hand, captures the update’s importance at the endpoint, determined by its relevance to the specific goal. Consequently, the magnitude of an update’s impact depends on the extent to which it satisfies the requirements of the targeted goal.
@@
We formulate a scheduling problem to maximize the expected discounted sum of the total grade of effectiveness of communicated updates, subject to a query cost constraint. In this context, we depart from classical EUT and leverage CPT to define long-term effectiveness. By solving this scheduling problem, we derive a class of effect-aware query scheduling policies for the hub. To this end, we propose two distinct approaches: (i) A model-based solution utilizing an iterative algorithm grounded in dynamic programming, and (ii) model-free solutions leveraging learning-based iterative algorithms, specifically leveraging a range of well-established deep reinforcement learning (DRL) methods. We evaluate the performance of our proposed solutions against benchmark approaches through simulations, focusing on their effectiveness. The results show that effect-aware scheduling significantly enhances both the long-term effectiveness and the system’s reliability, ensuring a minimum grade of effectiveness. Moreover, the findings reveal that effect-aware scheduling delivers substantial performance improvements in scenarios with stringent cost constraints. Additionally, the model-free approaches exhibit superior scalability compared to the model-based approach, making them more suitable for complex, real-world applications.
@@ I-A Related Works
Our work falls within the realm of query scheduling in pull-based communication systems. Although the pull-based model has been thoroughly explored, the scheduling problem aimed at enhancing the goal-oriented effectiveness of communicated updates under this model remains relatively unexplored. Existing studies on this problem, as highlighted in the literature, primarily focus on achieving accurate estimations of the source’s state over time [14, 15, 16, 17, 18]. These studies typically employ various forms of Kalman filters, relying on prior knowledge of the source’s dynamics at the estimator. The state of the source in these works evolves according to either linear or nonlinear dynamics, with the estimation error typically quantified using mean square error (MSE). The MSE is widely used as a fundamental metric for evaluating communication effectiveness and serves as the basis for formulating the scheduling problem in these approaches.
@@
The query scheduling problem we address is built on the grade of effectiveness metric, incorporating cumulative prospect theory for long-term effectiveness analysis. This framework provides a more comprehensive model that accounts for multiple attributes and integrates risk awareness. By enabling the evaluation of uncertain outcomes with subjective decision-making aspects, it is particularly suited for scenarios where attributes such as information freshness and usefulness must be considered simultaneously under risk conditions rather than optimizing a single attribute in isolation. In [28], we proposed a self-decision multiple access scheme to tackle a risk-agnostic form of the effectiveness enhancement problem. In this approach, sensing agents independently determine when to provide updates in response to queries, aiming to fulfill a specific goal. While the underlying intent is similar, the solution is designed from the perspective of the sensing agents. Moreover, the definition of the grade of effectiveness metric in this context differs, leading to a distinct objective function and a unique approach to addressing the problem.
@@ I-B Contributions
[…]
@@ II-A Scheduling Queries
The hub schedules queries about attributes to update $\mathbf{y}(t)$ in response to the needs of the AAs. In each time slot, the hub selects at most one attribute and queries an SA about that specific attribute. However, the hub operates without full knowledge of the evolution of the source. The query indicator for the $m$-th attribute from the $n$-th SA at the $t$-th slot is denoted by $\alpha_{nm}(t)=\{0,1\}$, where $\alpha_{nm}(t)=1$ indicates the query arrival, and $\alpha_{nm}(t)=0$ otherwise. Based on this, we can define $\mathbf{A}(t)=[\mathbf{a}_{1}(t),\dots,\mathbf{a}_{M}(t)]$ as the $N\times M$ query matrix, with each column $\mathbf{a}_{m}(t)=[\alpha_{1m}(t),\dots,\alpha_{Nm}(t)]^{T}$ representing an $N\times 1$ query vector corresponding on the $m$-th attribute. Therefore, the $m$-th element of $\mathbf{y}(t)$ is obtained as follows:
@@ II-B Grade of Effectiveness Definition
The grade of effectiveness (GoE) of an update on the $m$-th attribute is denoted by $\operatorname{GoE}_{m}(t)\in\mathbb{R}_{0}^{+}$ and is modeled as a composite function $f:\mathbb{R}_{0}^{+}\times\mathbb{R}_{0}^{+}\rightarrow\mathbb{R}_{0}^{+}$ of that update’s freshness and usefulness at the $t$-th slot. Thus, we can write
@@ III-A Problem Formulation
[…]
@@
The function $v_{\rm cpt}:\mathbb{R}_{0}^{+}\rightarrow\mathbb{R}$ represents the standard two-part CPT-based value function [12], which is concave for gains, convex for losses, and exhibits greater sensitivity to losses than to gains, reflecting loss aversion. The value function for gains (losses) is denoted by $v^{+}_{\rm cpt}(\cdot)$ ($v^{-}_{\rm cpt}(\cdot)$) and is monotonically non-decreasing (non-increasing) under positive (negative) prospects only. Otherwise, it evaluates to zero. Let $\operatorname{GoE}_{\rm ref}\in\mathbb{R}_{0}^{+}$ denote the reference point for the total GoE. In this context, any instance where $\operatorname{GoE}(t)\geq\operatorname{GoE}_{\rm ref},\forall t$, is considered a gain, while $\operatorname{GoE}(t)<\operatorname{GoE}_{\rm ref}$ refers to losses. For the induced query cost, we set the reference point to zero, making the prospect associated with the cost strictly positive.
@@ III-B CMDP Modeling
We can decompose the state space into $M+1$ disjoint subsets such that $\mathcal{S}=\bigcup_{m=0}^{M}\mathcal{S}_{m}$. For each $m=1,\dots,M$, the subset $\mathcal{S}_{m}$ corresponds to the set of states where $\Delta_{m}(t)<\Delta_{\rm max}$, while $\Delta_{m^{\prime}}(t)=\Delta_{\rm max},\forall{m^{\prime}}\neq m$, and $u_{m^{\prime}}(t)$ remains constant within this subset. Under the policy $\mathbf{1}_{N}^{T}\mathbf{a}_{m}(t)=1$, the subset $\mathcal{S}_{m}$ indicates the set of recurrent states. The subset $\mathcal{S}_{0}$ consists of the remaining states, all of which are transient under some policy. Therefore, the weak accessibility condition holds [29, Definition 4.2.2], as the process can reach any recurrent state from a transient one with a proper policy. ∎
@@
With the weak accessibility condition satisfied for the defined CMDP model, the expected discounted sum of the CPT-based total GoE in $\mathcal{P}$ remains invariant across all initial states [29, Proposition 4.2.3]. Consequently, $v_{\rm cpt}\big{(}\!\operatorname{GoE}(t)\big{)}$, is independent of $\operatorname{GoE}(0)$, for all $t\geq 1$.
@@
Satisfying the weak accessibility condition guarantees the existence of a class of stationary optimal policies, denoted as $\pi^{*}$, for $\mathcal{P}$ in (III-A), which are unichain [29, Proposition 4.2.6].
@@ IV-A Dual Problem
[…]
@@
There exists a saddle point that ensures the convergence of the original scheduling problem, i.e., $\mathcal{P}$, and the dual one, i.e., $\widehat{\mathcal{P}}$, to the same values. This is due to (i) the finite state space of the CMDP model, satisfying the growth condition [30], and (ii) the boundness of the immediate rewards in that model. In this sense, the following equality holds [30, Corollary 12.2]:
[…]
where $\mu^{*}$ denotes a non-negative optimal Lagrange multiplier for some $\pi$, which is feasible owing to (i) and (ii) under the Slater’s condition [30, Theorem 12.8].
@@ IV-B Iterative Algorithm
Input: Given parameters $M$, $N$, $T\gg 0$, and $\gamma$, $C_{\rm max}$. CMDP’s state space, i.e., $\mathcal{S}$, and action space, i.e., $\mathcal{A}$. Shapes of $v_{\rm cpt}(\cdot)$, $w_{\rm cpt}(\cdot)$, and $f_{c}(\cdot)$. Tolerance $\epsilon_{\mu}$. Mixing factor $\eta$. Initial values $l\leftarrow 0$, $\mu^{(0)}\leftarrow 0$, $\mu_{-}^{(0)}\leftarrow 0$, $\mu_{+}^{(0)}>0$, $\pi^{-}\leftarrow 0$, and $\pi^{+}\leftarrow 0$.
[…]
@@
[…]
@@
[…]
@@
20            step up $i\leftarrow i+1$, and goto 1.
@@ IV-B1 Computing $\pi^{*}$
using the Bellman equation [31], where $Q^{(i)}_{\pi}(s,a)$ denotes the state-action value function (Q-function) defined as follows:
@@
based on the span seminorm [32, Section 6.6.1]. Given Corollary 2 and the fact that every optimal class of policies has an aperiodic transition matrix under the modeled CMDP, the criterion in (14) is guaranteed to be satisfied after a finite number of iterations [32, Theorem 8.5.4].
@@ IV-B2 Computing $\mu^{*}$
The search starts from an initial interval $[\mu_{-}^{(0)},\mu_{+}^{(0)}]$ complying with $\mathcal{L}(\pi^{*},\mu_{-}^{(0)})\mathcal{L}(\pi^{*},\mu_{+}^{(0)})<0$ given $\pi^{*}$ from the inner loop. In each subsequent step, the Lagrange multiplier is updated to the midpoint of the current interval. In this context, we have $\mu^{(l)}=\frac{\mu_{-}^{(l)}+\mu_{+}^{(l)}}{2}$ for the $l$-th step, where $l\in\mathbb{N}$. The interval is updated after computing $\pi^{*}$ based on the new multiplier and evaluating whether the cost constraint is met. The bisection search terminates when the stopping criterion, as defined below, is met:
@@
Thus, the bisection search convergences to the optimal Lagrange multiplier within a finite number of steps [33].
@@ V-A Model-Free Environment
To address the limitations of relying on a specific model, we consider an environment that encompasses the source, sensing agents, error-prone channels for transporting noisy updates to the hub, and the evolving knowledge base. In this mode-free approach, the hub interacts with the environment by performing actions, making observations, and receiving rewards, all without prior knowledge of the source’s dynamics, the potential usefulness of the updates to be queried, or the status of the update channels.

### v2

@@ I Introduction
When analyzing long-term effectiveness, the axioms of classical expected utility theory (EUT), where the weight of an occurrence is determined by its probability, could be challenged in the context of risk-aware semantic decision-making. Decisions involving risky prospects, perceptual context evaluation, and semantic utility valuation reveal several effects that the utility theory fails to capture. One prominent effect is certainty, which suggests that merely probable outcomes are undervalued compared to certain ones. For example, in human-centric applications, there is often a preference to guarantee achieving a significant portion of a goal rather than risk achieving the entire goal with less certainty. Decision-making involving human agents and human-in-the-loop control of critical applications tends to be more conservative, prioritizing safety and performance guarantees over risk-taking, even when the latter has the potential to maximize the utility function. Cumulative prospect theory (CPT) was developed as an alternative framework to EUT, addressing subjectivity and behavioral effects in decision-making [10, 11, 12]. This theory offers a more refined and nuanced framework that captures how individuals perceive gains and losses relative to a reference point. It maps the outcomes to utility via a value function and transforms the perceived probabilities of those outcomes into decision weights using a weighting function. These elements allow CPT to account for behavioral phenomena such as loss aversion and the overweighting of rare events. This makes CPT especially relevant in scenarios where infrequent but impactful updates matter, such as semantic communications, which prioritize information based on perceived significance rather than mere frequency. Consequently, CPT provides a more realistic basis for modeling decision-making under uncertainty.
@@
In this paper, we investigate a pull-based end-to-end status update system in which multiple sensing agents monitor a source and provide updates to a group of actuation agents, each taking actions and working toward achieving heterogeneous goals at the endpoint. The source comprises multiple attributes, and updates about these attributes are routed through a hub that dynamically determines which specific attribute to query at any given time. To assess the effectiveness of updates, we utilize the grade of effectiveness concept introduced in [13], incorporating two key semantic metrics: freshness and usefulness. Freshness quantifies how quickly an update becomes outdated after successfully reaching its endpoint. Usefulness, on the other hand, captures the update’s importance at the endpoint, determined by its relevance to the specific goal. Consequently, the magnitude of an update’s impact depends on the extent to which it satisfies the requirements of the goal.
@@
We formulate a scheduling problem to maximize the expected discounted sum of the total grade of effectiveness of communicated updates, subject to a query cost constraint. In this context, we depart from the classical EUT and leverage CPT to define long-term effectiveness. By solving this scheduling problem, we derive a class of effect-aware query scheduling policies for the hub. To this end, we propose two distinct approaches: (i) A model-based solution utilizing an iterative algorithm grounded in dynamic programming, and (ii) model-free solutions leveraging learning-based iterative algorithms, specifically using a range of well-established deep reinforcement learning (DRL) methods. We evaluate the performance of our proposed solutions against benchmark approaches through simulations, focusing on their effectiveness. The results show that effect-aware scheduling significantly enhances both the long-term effectiveness and the system’s reliability, ensuring a minimum grade of effectiveness. In addition, the findings reveal that effect-aware scheduling delivers substantial performance improvements in scenarios with stringent cost constraints. Additionally, model-free approaches exhibit superior scalability compared to the model-based approach, making them more suitable for complex, real-world applications.
@@ I-A Related Works
Our work falls within the realm of query scheduling in pull-based communication systems. Although the pull-based model has been thoroughly explored, the scheduling problem aimed at enhancing the goal-oriented effectiveness of communicated updates under this model remains relatively unexplored. Existing studies on this problem, as highlighted in the literature, focus mainly on achieving accurate estimates of the source’s state over time [14, 15, 16, 17, 18]. These studies typically employ various forms of Kalman filters, relying on prior knowledge of the source’s dynamics at the estimator. The state of the source in these works evolves according to linear or nonlinear dynamics, with the estimation error typically quantified using mean square error (MSE). The MSE is widely used as a fundamental metric for evaluating communication effectiveness and serves as the basis for formulating the scheduling problem in these approaches.
@@
The query scheduling problem we address is built on the grade of effectiveness metric, incorporating cumulative prospect theory for long-term effectiveness analysis. This framework provides a more comprehensive model that accounts for multiple attributes and integrates risk awareness. By enabling the evaluation of uncertain outcomes with subjective decision-making aspects, it is particularly suited for scenarios where attributes such as information freshness and usefulness must be considered simultaneously under risk conditions rather than optimizing a single attribute in isolation. In [28], we proposed a self-decision multiple access scheme to address a risk-agnostic variant of the effectiveness enhancement problem. In that framework, sensing agents autonomously decide whether and when to provide updates in response to incoming queries, guided by a predefined objective. While the overarching goal aligns with the one pursued here, the approach in [28] is decentralized and agent-centric, in contrast to the centralized strategy proposed in this work. As such, the two methodologies address complementary layers of the decision-making process, offering a more holistic perspective when considered together. Furthermore, the form of the grade of effectiveness metric in this context differs, leading to a distinct objective function and a unique approach to address the problem.
@@ I-B Contributions
Notations: $\mathbb{R}$, $\mathbb{R}_{0}^{+}$, and $\mathbb{N}$ represent the sets of real, non-negative real, and natural numbers, respectively. $\mathbf{1}_{x}$ denotes an all-ones column vector of size $x$. $\mathbb{E}[\cdot]$ is the expectation operator, $\lvert\cdot\rvert$ indicates the absolute value operator, $\lceil\cdot\rceil$ is the ceiling operator, and $\mathcal{O}(\cdot)$ describes a function’s growth rate.
@@ II-A Scheduling Queries
The hub schedules queries about attributes to update $\mathbf{y}(t)$ in response to the needs of the AAs. In each time slot, the hub selects at most one [^2] This condition simplifies the presentation of the core idea, ensuring clarity and focus while minimizing unnecessary complexity. attribute and queries an SA about that specific attribute. However, the hub operates without full knowledge of the evolution of the source. The query indicator for the $m$-th attribute from the $n$-th SA at the $t$-th slot is denoted by $\alpha_{nm}(t)=\{0,1\}$, where $\alpha_{nm}(t)=1$ indicates the query arrival, and $\alpha_{nm}(t)=0$ otherwise. Based on this, we can define $\mathbf{A}(t)=[\mathbf{a}_{1}(t),\dots,\mathbf{a}_{M}(t)]$ as the $N\times M$ query matrix, with each column $\mathbf{a}_{m}(t)=[\alpha_{1m}(t),\dots,\alpha_{Nm}(t)]^{T}$ representing an $N\times 1$ query vector corresponding on the $m$-th attribute. Therefore, the $m$-th element of $\mathbf{y}(t)$ is obtained as follows:
@@ II-B Grade of Effectiveness Definition
The grade of effectiveness (GoE) of an update on the $m$-th attribute is denoted by $\operatorname{GoE}_{m}(t)\in\mathbb{R}_{0}^{+}$ and is modeled as a composite function $f:\mathbb{R}_{0}^{+}\times\mathbb{R}_{0}^{+}\rightarrow\mathbb{R}_{0}^{+}$ of that update’s freshness and usefulness at the $t$-th slot. Inspired from [29, Section III] and [13, Section III-A], we can write
@@ III-A Problem Formulation
$\mathcal{P}:$~ $\underset{\pi}{\operatorname{max}}~~\mathbb{E}_{\pi}\!\left[\,\sum_{t=0}^{\infty}\gamma^{t}v_{\rm cpt}\big{(}\!\operatorname{GoE}(t)\big{)}\,\Big{|}\operatorname{GoE}(0),\pi\right]$
@@
[…]
@@ III-B CMDP Modeling
We can decompose the state space into $M+1$ disjoint subsets such that $\mathcal{S}=\bigcup_{m=0}^{M}\mathcal{S}_{m}$. For each $m=1,\dots,M$, the subset $\mathcal{S}_{m}$ corresponds to the set of states where $\Delta_{m}(t)<\Delta_{\rm max}$, while $\Delta_{m^{\prime}}(t)=\Delta_{\rm max},\forall{m^{\prime}}\neq m$, and $u_{m^{\prime}}(t)$ remains constant within this subset. Under the policy $\mathbf{1}_{N}^{T}\mathbf{a}_{m}(t)=1$, the subset $\mathcal{S}_{m}$ indicates the set of recurrent states. The subset $\mathcal{S}_{0}$ consists of the remaining states, all of which are transient under some policy. Therefore, the weak accessibility condition holds [30, Definition 4.2.2], as the process can reach any recurrent state from a transient one with a proper policy. ∎
@@
With the weak accessibility condition satisfied for the defined CMDP model, the expected discounted sum of the CPT-based total GoE in $\mathcal{P}$ remains invariant across all initial states [30, Proposition 4.2.3]. Consequently, $v_{\rm cpt}\big{(}\!\operatorname{GoE}(t)\big{)}$, is independent of $\operatorname{GoE}(0)$, for all $t\geq 1$.
@@
Satisfying the weak accessibility condition guarantees the existence of a class of stationary optimal policies, denoted as $\pi^{*}$, for $\mathcal{P}$ in (III-A), which are unichain [30, Proposition 4.2.6].
@@ IV-A Dual Problem
$\widehat{\mathcal{P}}$:~ $\underset{\mu}{\operatorname{inf}}~\underset{\pi}{\operatorname{sup}}~~\mathcal{L}(\pi,\mu)$
@@
There exists a saddle point that ensures the convergence of the original scheduling problem, i.e., $\mathcal{P}$, and the dual one, i.e., $\widehat{\mathcal{P}}$, to the same values. This is due to (i) the finite state space of the CMDP model, satisfying the growth condition [31], and (ii) the boundness of the immediate rewards in that model. In this sense, the following equality holds [31, Corollary 12.2]:
$\underset{\mu\geq 0}{\operatorname{inf}}~\underset{\pi}{\operatorname{sup}}~\mathcal{L}(\pi,\mu)=\underset{\mu\geq 0}{\operatorname{inf}}~\mathcal{L}(\pi^{*},\mu)=\underset{\pi}{\operatorname{sup}}~\mathcal{L}(\pi,\mu^{*})$ (10)
where $\mu^{*}$ denotes a non-negative optimal Lagrange multiplier for some $\pi$, which is feasible owing to (i) and (ii) under the Slater’s condition [31, Theorem 12.8].
@@ IV-B Iterative Algorithm
Input: Given parameters $M$, NN, $T\gg 0$, and $\gamma$, $C_{\rm max}$. CMDP’s state space, i.e., $\mathcal{S}$, and action space, i.e., $\mathcal{A}$. Shapes of $v_{\rm cpt}(\cdot)$, $w_{\rm cpt}(\cdot)$, and $f_{c}(\cdot)$. Tolerance $\epsilon_{\mu}$. Mixing factor $\eta$. Initial values $l\leftarrow 0$, $\mu^{(0)}\leftarrow 0$, $\mu_{-}^{(0)}\leftarrow 0$, $\mu_{+}^{(0)}>0$, $\pi^{-}\leftarrow 0$, and $\pi^{+}\leftarrow 0$.
1 Initialize $\pi^{*}(s),\forall s\in\mathcal{S}$, via $\text{derive}\_\text{policy}$($\mu^{(0)}$).
@@
7  Improve $\pi^{*}\leftarrow$ $\text{derive}\_\text{policy}$($\mu^{(l)}$).
@@
9    $\mu_{-}^{(l)}\leftarrow\mu^{(l)}$, and $\pi^{-}\leftarrow$ $\text{derive}\_\text{policy}$($\mu_{-}^{(l)}$).
10 else $\mu_{+}^{(l)}\leftarrow\mu^{(l)}$, and $\pi^{+}\leftarrow$ $\text{derive}\_\text{policy}$($\mu_{+}^{(l)}$).
@@
20    increment $i\leftarrow i+1$, and goto 1.
@@ IV-B1 Computing $\pi^{*}$
using the Bellman equation [32], where $Q^{(i)}_{\pi}(s,a)$ denotes the state–action value function (Q-function) defined as follows:
@@
based on the span seminorm [33, Section 6.6.1]. Given Corollary 2 and the fact that every optimal class of policies has an aperiodic transition matrix under the modeled CMDP, the criterion in (14) is guaranteed to be satisfied after a finite number of iterations [33, Theorem 8.5.4].
@@ IV-B2 Computing $\mu^{*}$
The search starts from an initial interval $\mu_{-}^{(0)},\mu_{+}^{(0)}]$ complying with $\mathcal{L}(\pi^{*},\mu_{-}^{(0)})\mathcal{L}(\pi^{*},\mu_{+}^{(0)})<0$ given $\pi^{*}$ from the inner loop. In each subsequent step, the Lagrange multiplier is updated to the midpoint of the current interval. In this context, we have $\mu^{(l)}=\frac{\mu_{-}^{(l)}+\mu_{+}^{(l)}}{2}$ for the $l$-th step, where $l\in\mathbb{N}$. The interval is updated after computing $\pi^{*}$ based on the new multiplier and evaluating whether the cost constraint is met. The bisection search terminates when the stopping criterion, as defined below, is met:
@@
Thus, the bisection search converges to the optimal Lagrange multiplier within a finite number of steps [34].
@@ V-A Model-Free Environment
To address the limitations of relying on a specific model, we consider an environment that encompasses the source, sensing agents, error-prone channels for transporting noisy updates to the hub, and the evolving knowledge base. In this model-free approach, the hub interacts with the environment by performing actions, making observations, and receiving rewards, all without prior knowledge of the source’s dynamics, the potential usefulness of the updates to be queried, or the status of the update channels.